### v1

# Evaluating Test-Time Scaling LLMs for Legal Reasoning: OpenAI o1, DeepSeek-R1, and Beyond

Abstract
Recently, Test-Time Scaling Large Language Models (LLMs), such as DeepSeek-R1 and OpenAI o1, have demonstrated exceptional capabilities across various domains and tasks, particularly in reasoning. While these models have shown impressive performance on general language tasks, their effectiveness in specialized fields like legal remains unclear. To address this, we present a preliminary evaluation of LLMs in various legal scenarios, covering both Chinese and English legal tasks. Our analysis includes 9 LLMs and 17 legal tasks, with a focus on newly published and more complex challenges such as multi-defendant legal judgments and legal argument reasoning. Our findings indicate that, despite DeepSeek-R1 and OpenAI o1 being among the most powerful models, their legal reasoning capabilities are still lacking. Specifically, these models score below 80% on seven Chinese legal reasoning tasks and below 80% on two English legal reasoning tasks. This suggests that, even among the most advanced reasoning models, legal reasoning abilities remain underdeveloped.
1 Introduction
Recent advancements in large language models (LLMs) have markedly reduced the gap toward Artificial General Intelligence (AGI)  Meta (2024); Team (2024); Openai (2024a, b); team (2025); Anthropic (2025). Exemplary models such as OpenAI GPT-4o  Openai (2024a), Qwen2.5  Team (2024), and Meta Llama-3.1  Meta (2024) highlight this progress. Among them, OpenAI has introduced a series of models, referred as o1  Openai (2024b), which improve the reasoning performance across diverse domains including mathematics, coding, and scientific inquiry by increasing the test-time computation through extending the length of the chain-of-thought. However, o1 are proprietary models, and their technical details and weights have not been disclosed. Following OpenAI o1, several studies have been explored to match their comprehensive reasoning capabilities, such as the Process Reward Model (PRM)  Lightman et al. (2023); Uesato et al. (2022); Wang et al. (2024), Reinforcement Learning(RL) techniques  Kumar et al. (2024), and search algorithms like Monte Carlo Tree Search and Beam Search  Feng et al. (2024); Trinh et al. (2024). However, the performance is still lacking.
Figure 1: Overall results of the LLMs in Chinese and English legal tasks. We show performance on all legal datasets, excluding Sara_N for comparative analysis. DeepSeek-R1 and o1-preview demonstrate a clear performance advantage over the General and other TTS LLMs.
More recently, DeepSeek has open-sourced its test-time compute LLM, DeepSeek-R1 DeepSeek-AI et al. (2025). DeepSeek-AI et al. (2025) found that large-scale reinforcement learning (RL) with simple rule-based reward can already incentive the reasoning ability of LLMs, e.g., the occurrence of an "aha moment", where the model is referred to as Deekseek-R1-Zero. However, only training with RL leads to poor readability and language mixing issues. To address these challenges, a four-stage pipeline is proposed, comprising: (i) a cold-start stage, (ii) reasoning-oriented reinforcement learning, (iii) rejection sampling and supervised fine-tuning, and (iv) reinforcement learning across all scenarios. Upon completing these stages, the resulting model, DeepSeek-R1, effectively mitigates the limitations of pure reinforcement learning while demonstrating strong reasoning capabilities and competitive performance across a broad spectrum of general-purpose tasks. Notably, DeepSeek-R1 outperforms OpenAI o1-1217 on several reasoning benchmarks, attaining an accuracy of 79.8% and 97.3% on AIME 2024 and MATH-500, respectively. Concurrently, other test-time scaling LLMs have emerged, including Qwen Team’s QWQ-32B-Preview  Qwen Team (2024) and Zhipu AI’s GLM-zero-preview  [^1] https://bigmodel.cn/dev/api/normal-model/glm-zero-preview.
Despite significant progress in enhancing LLMs’ reasoning capabilities, questions remain about their performance in legal reasoning tasks, which require not only relevant legal knowledge but also the ability to apply that knowledge through reasoning. Do models that perform well in general reasoning exhibit comparable proficiency in the legal domain? If not, what are the underlying reasons for this discrepancy? Prior research has explored GPT-4’s performance across various legal tasks, including comparisons with earlier GPT models on legal text annotation  Savelka and Ashley (2023), generating factual explanations of legislative terms  Savelka et al. (2023), and conducting thematic analysis in empirical legal studies  Drápal et al. (2023). However, none of these studies offer a comprehensive and in-depth analysis of LLMs’ legal reasoning capabilities.
To address this gap, we conduct a comprehensive evaluation of 17 legal reasoning tasks (7 in English and 10 in Chinese) from prior research, testing a range of representative LLMs. Our assessment includes both test-time scaling LLMs and general-purpose LLMs, aiming to systematically analyze their performance in legal reasoning tasks. Key findings from our evaluation include:
  1. 1.
    In the Chinese legal reasoning tasks, the DeepSeek-R1 model demonstrates the highest performance among the evaluated models, with OpenAI o1-preview following closely behind. However, all evaluated LLMs encounter significant challenges when dealing with tasks that demand complex reasoning capabilities, such as controversial focus extraction, judicial examination, and multi-defendant legal judgment prediction.
  2. 2.
    In English legal reasoning tasks, all evaluated LLMs generally exhibit stronger performance compared to their Chinese task. Notably, DeepSeek-R1 and OpenAI o1-preview perform comparably, with each achieving the highest score in three tasks. However, even these advanced LLMs struggle with tasks related to judicial ethics and complex tax calculations.
  3. 3.
    Across all legal tasks, the performance of different TTS LLMs models varies significantly. Notably, o1-preview and DeepSeek-R1 outperform other models such as QWQ-32B-Preview and GLM-zero-preview. In the future, further refinement and optimization of these models may be necessary.
  4. 4.
    A detailed error analysis of DeepSeek-R1’s performance on several representative Chinese and English legal tasks, reveals that the model has shortcomings in aspects such as lacking sufficient and updated legal knowledge, struggling to understand legal issues, and factual hallucinations in its thinking process, which offers new insights and directions for enhancing LLMs’ legal reasoning skills.
2 Related Work
2.1 Test-Time Scaling LLMs
Recently, Test-Time Scaling(TTS) has been an effective way to improve the performance of LLMs. Unlike traditional optimization methods, TTS focuses on optimizing the model’s output during the inference phase through additional computational resources, without requiring modifications to the model’s parameters or structure. This strategy aims to enhance the accuracy and efficiency of the model by dynamically adjusting the reasoning process, particularly when dealing with complex tasks, improving the model’s performance.
A number of works have proposed different methods for enabling LLMs to use test-time scaling for reasoning  Snell et al. (2024); Liu et al. (2025). Existing related works mainly focus on areas such as verifier optimization, proposal distribution refinement, and search algorithms. The verifier optimization involves using process-based verifiers, such as process reward models (PRM)  Lightman et al. (2023); Uesato et al. (2022); Wang et al. (2024) to incrementally evaluate the reasoning steps, which effectively improves the model’s performance in complex reasoning tasks. Additionally, proposal distribution refinement helps LLMs generate more accurate and refined candidate answers during the test phase. This strategy of adjusting the proposal distribution further enhances the model’s reasoning ability, allowing it to iteratively refine its initial outputs and gradually approach the optimal solution. Examples of such methods include STaR and ReSTEM  Singh et al. (2024); Zelikman et al. (2022). These techniques do not require any additional input tokens but instead fine-tune the model to induce an improved proposal distribution. Alternatively, techniques like self-critique  Bai et al. (2022); Du et al. (2023); Madaan et al. (2023); Saunders et al. (2022) enable the model to improve its proposal distribution during test time by instructing it to critique and revise its outputs iteratively. In terms of search algorithms, integrating techniques like Beam Search and Monte Carlo Tree Search (MCTS)  Feng et al. (2024); Trinh et al. (2024) allows for more refined exploration of solution paths, which further enhances the accuracy of reasoning results. A prime example of test-time scaling LLMs is OpenAI’s o1 series models  Openai (2024b), which enhance reasoning capabilities by extending the Chain-of-Thought (CoT) process during the test-time. Besides, Alibaba’s QwQ-32B-Preview  Qwen Team (2024) model showcases the application of test-time computation in mathematical problem-solving. Additionally, the GLM-zero-preview  [^2] https://bigmodel.cn/dev/api/normal-model/glm-zero-preview released by Zhipu AI is another notable example. Recently, by exploring how LLMs can develop reasoning capabilities without supervised data, DeepSeek focuses on the potential for self-evolution through pure RL processes and successfully develops the DeepSeek-R1 model  DeepSeek-AI et al. (2025). Despite having far fewer parameters than o1, DeepSeek-R1 performs on par with OpenAI’s o1-1217 model. Furthermore, DeepSeek-R1 has demonstrated exceptional performance in coding tasks, achieving a 2029 Elo rating on Codeforces, outclassing 96.3% of participants.
However, amid the excitement, there are also concerns, as the TTS LLMs may generate harmful content. If errors occur in legal applications, the consequences can be severe  Weiser (2023). Therefore, it is urgent to test the performance of LLMs in legal contexts. Therefore, this paper aims to explore whether the enhanced abilities of the TTS LLMs effectively transfer to the legal domain and improve their performance on legal reasoning tasks.
2.2 Legal Reasoning Benchmarks
Understanding the capabilities of LLMs in legal tasks, particularly legal reasoning, is a key focus of research  Blair-Stanek et al. (2023); Trozze et al. (2024). Legal reasoning ability refers to whether LLMs can effectively leverage their legal knowledge and perform the necessary reasoning steps to solve real-world legal tasks  Li et al. (2024a). This requires not only relevant legal knowledge but also the ability to apply that knowledge through reasoning. For instance, in Chinese judicial applications, tasks like legal document generation  Iu and Wong (2023), legal case retrieval  Ma et al. (2021), legal question answering  Hu et al. (2025), and judgment prediction  Gan et al. (2021, 2022); Jiang and Yang (2023); Wei et al. (2025) require that LLMs process precise legal knowledge and complex legal contexts. These tasks often involve specialized components such as judicial interpretation and reasoning.
Many researchers propose various legal reasoning tasks and datasets to assess LLMs’ performance in legal reasoning, covering different legal systems like the civil law system and the common law system  Magesh et al. (2024).  Chlapanis et al. (2024) introduce Legal Argument Reasoning (LAR), a novel task designed to evaluate the legal reasoning capabilities of LLMs.  Deng et al. (2023) develop a syllogistic reasoning dataset for legal judgment analysis, manually curating 11,239 criminal cases that cover 4 criminal elements, 80 charges, and 124 articles.  Joshi et al. (2024) propose the Indian Legal Text Understanding and Reasoning Benchmark (IL-TUR), which includes monolingual (English, Hindi) and multilingual (9 Indian languages) tasks to assess the understanding and reasoning capabilities of Indian legal documents. In addition,  Kang et al. (2023) create a novel corpus focused on Contract Acts in Malaysia and the Australian Social Act for Dependent Children, where ChatGPT is applied using the IRAC method to perform legal analysis.
Several legal task benchmarks are proposed to evaluate whether LLMs meet the high standards required for legal practice, such as Laiw  Dai et al. (2025). In the common law system, LegalBench  Guha et al. (2023) is a widely used legal reasoning benchmark, collaboratively built to consist of 162 tasks across six types of legal reasoning. In the civil law system,  Fei et al. (2024) introduce LawBench, a comprehensive evaluation benchmark that precisely assesses LLMs’ legal capabilities at three cognitive levels: (1) Legal knowledge memorization, (2) Legal knowledge understanding, and (3) Legal knowledge application. Moreover,  Li et al. (2024a) introduce LexEval, a standardized Chinese legal benchmark that focuses not only on LLMs’ ability to apply fundamental legal knowledge but also on examining ethical issues in their application.
However, legal systems vary greatly across countries. The Chinese legal system, rooted in the civil law tradition, differs significantly from the common law systems widely used in the U.S. and the U.K. Therefore, this paper selects a set of existing Chinese and English legal task datasets, along with newly proposed datasets, to comprehensively evaluate the diverse capabilities of LLMs.
3 Evaluation Setting
To comprehensively evaluate the legal reasoning capabilities of LLMs, we compile a diverse set of legal reasoning tasks in both Chinese and English.
3.1 Legal Reasoning Task
In this section, we introduce the legal reasoning tasks we have compiled, which include ten Chinese legal tasks from eleven distinct datasets, along with seven English legal tasks.
3.1.1 Chinese Legal Tasks
Table 1: Chinese Legal Tasks
Task Dataset Source Metric Test Size
Legal Calculation(LC) LC LexEval Acc 234
Legal Multi-hop Reasoning(LMHR) LMHR LexEval Acc 200
Legal Judgment Prediction(LJP) CAIL2018 CAIL2018 F1 300
Multi-Defendant Legal Judgment Prediction(MDLJP) CMDL  Huang et al. (2024) F1 300
Multi-Defendant Legal Judgment Prediction(MDLJP) MultiLJP  Lyu et al. (2023) F1 300
Multi-Defendant Charge Prediction(MDCP) MUD  Wei et al. (2024) F1 175
Multi-segment Legal Reading Comprehension(MSLRC) MSLRC CAIL2021 F1 200
Controversial Focus Extraction(CFE) CFE LAIC2021 F1 200
Interactive Argument-Pair Extraction(IAPE) ArgMine CAIL2023 Acc 200
Article Recitation(AR) AR LawBench Rouge-L 200
Judicial Examination(JE) JE JEC-QA Acc 300
As shown in table 1, we provide the source and test size for each dataset associated with the tasks. The detailed descriptions of the datasets and tasks are as follows:
Legal Calculation: The legal calculation task involves answering multiple-choice questions that require legal computations. For each question, the model must select the single correct option from A, B, C, or D. This task is evaluated on the LC dataset derived from LexEval  Li et al. (2024b), a comprehensive Chinese legal benchmark for assessing large language models, using accuracy as the evaluation metric.
Legal Multi-hop Reasoning: This task assesses the legal knowledge and reasoning capabilities of LLMs. The input consists of multiple-choice questions related to legal matters, and the model’s output is the correct answer(s) from the provided options, which may include one or more correct choices. The LMHR dataset, sourced from LexEval, is used for this task, with accuracy as the evaluation metric.
Legal Judgment Prediction: This task focuses on legal judgment prediction for single-defendant cases based on the CAIL2018 dataset. The input includes a detailed description of case facts and defendant information, while the output provides judgment results for three subtasks: charge prediction, article prediction, and sentence prediction. The evaluation employs the same metrics as those used in CAIL2024  [^3] https://github.com/china-ai-law-challenge/CAIL2024/tree/main/drdz, with the calculations detailed as follows:
For a given case $c$ with $n$ defendants, consider a defendant $d$ who is charged with $m_{1}$ crimes. If the model predicts $m_{2}$ crimes for this defendant, with $m_{3}$ of them being correct, the precision ($P$), recall ($R$), and F1 score ($F1$) for the charge and article prediction subtasks for this defendant are defined as follows:
$P_{d}^{c}=\frac{m_{3}}{m_{2}},R_{d}^{c}=\frac{m_{3}}{m_{1}},F1_{d}^{c}=\frac{2\cdot P_{d}^{c}\cdot R_{d}^{c}}{P_{d}^{c}+R_{d}^{c}}$ (1)
The P, R, and F1 Score for this case are calculated as follows:
$P_{c}=\frac{\sum_{i=1}^{n}P_{i}^{c}}{n}$ (2)
$R_{c}=\frac{\sum_{i=1}^{n}R_{i}^{c}}{n}$ (3)
$F1_{c}=\frac{\sum_{i=1}^{n}F1_{i}^{c}}{n}$ (4)
For the entire dataset, these metrics are weighted by $w_{c}=\log_{2}n$:
$P=\frac{\sum w_{c}P_{c}}{\sum w_{c}},R=\frac{\sum w_{c}R_{c}}{\sum w_{c}},F1=\frac{\sum w_{c}F1_{c}}{\sum w_{c}}$ (5)
The metric of sentence prediction for case $c$ is evaluated using the accuracy metric. For a given case $c$ with $n$ defendants, if $k$ defendants have correctly predicted sentences, then:
$Acc_{c}=\frac{k}{n}$ (6)
The sentence accuracy for the entire dataset is:
$Acc=\frac{\sum w_{c}Acc_{c}}{\sum w_{c}},\quad w_{c}=\log_{2}n$ (7)
Finally, the overall F1 score, combining the metrics for the three subtasks, is calculated as:
[…]
Multi-Defendant Legal Judgment Prediction: This task focuses on predicting legal judgments in cases involving multiple defendants. The task utilizes two datasets: CMDL from  Huang et al. (2024) and MultiLJP from  Lyu et al. (2023), and employs the same evaluation metrics as used in the LJP task.
Multi-Defendant Charge Prediction: This task focuses on predicting charges for multiple defendants. Given the case facts as input, the goal is to determine the charges committed by each defendant. The dataset used is MUD from  Wei et al. (2024), and the evaluation metric is analogous to that of the charge prediction subtask in the LJP task.
Multi-segment Legal Reading Comprehension: This task involves multi-segment questions, where the answers are derived by extracting and combining multiple segments from the legal text. The dataset employed is MSLRC from CAIL2021. To evaluate LLMs performance on this task, we designed a metric tailored to its characteristics. Here, both the ground truth $G=\{g_{1},g_{2},\dots,g_{n}\}$ and the model output $E=\{e_{1},e_{2},\dots,e_{m}\}$ are lists of legal elements that answer the question within its legal context. A pre-trained language model is used to automatically assess the semantic similarity between the elements in $G$ and $E$. Finally, the F1 score is computed as the evaluation metric, calculated as follows:
$P=\frac{N}{m},R=\frac{N}{n},F1=\frac{2PR}{P+R}$ (9)
where $N$ represents the number of correctly predicted legal elements in $E$, $m$ is the total number of elements in the output $E$, and $n$ is the total number of elements in the ground truth $G$.
Controversial Focus Extraction: This task entails identifying dispute issues based on the claims and defenses from both the plaintiff and defendant. The output is a list of controversial focus indices extracted from the case facts. LLMs performance is assessed using the F1 score, calculated similarly to the MSLRC task. However, rather than relying on a pre-trained language model for semantic interpretation, we directly verify whether the predicted indices match the ground truth indices.
Interactive Argument-Pair Extraction: This task aims to extract interaction argument pairs by identifying the defense counter-argument that corresponds to a given plaintiff’s argument. The input comprises the plaintiff’s argument along with five candidate defense arguments, and the output is the selected counter-argument. Performance is measured using accuracy.
Article Recitation: This task assesses LLMs’ ability to recall legal knowledge by prompting them to recite the content of legal articles based on their reference numbers. It examines their proficiency in memorizing key legal concepts, terminology, and provisions. The dataset is sourced from the comprehensive LawBench evaluation benchmark  Fei et al. (2024), and Rouge-L is employed as the evaluation metric.
Judicial Examination: This task requires LLMs to output the final answers to the questions from JEC-QA  Zhong et al. (2019), which is the largest question answering dataset in the legal domain, collected from the National Judicial Examination of China. We randomly test the 300 cases from the concept comprehension questions and scenario analysis questions, which require the ability of logical reasoning. The performance of LLMs is measured using accuracy.
3.1.2 English Legal Tasks
Table 2: English Legal Tasks
Task and Dataset Source Metric Test Size
Legal Reasoning Causality(LRC) LegalBench Acc 55
Citation Prediction Classification(CPC) LegalBench Acc 108
NYS Judicial Ethics(NYSJE) LegalBench Acc 292
Sara Numeric(Sara_N) LegalBench Mse 96
Sara Entailment(Sara_E) LegalBench Acc 272
Supreme Court Assessment of Legal Reasoning(Scalr) LegalBench Acc 172
Legal Argument Reasoning(LAR)  Chlapanis et al. (2024) Acc 200
The English legal reasoning tasks are mainly sourced from LegalBench  Guha et al. (2023), a collaboratively constructed legal reasoning benchmark consisting of 162 tasks covering six different types of legal reasoning. Besides, we also add a new legal argument reasoning task proposed by  Chlapanis et al. (2024). The tasks are listed as follows:
Legal Reasoning Causality: This task aims to classify whether an excerpt from a district court opinion relies on statistical evidence in its reasoning.
Citation Prediction Classification: The task requires determining whether a given case citation supports a legal statement, based on the provided legal statement and citation.
NYS Judicial Ethics: In this task, LLMs are required to determine whether a question violates judicial ethics in the New York State Unified Court System. The dataset consists of real ethical scenarios, reformulated into questions to evaluate the models’ understanding of ethical rules and their application in different judicial contexts.
Sara Numeric: In this task, the LLMs should determine how much tax an individual owes given a statute and accompanying facts. The dataset in this task is from the StAtutory Reasoning Assessment(SARA), it contains a set of statutes and summaries of facts paired with a numerical question. Additionally, we use Mean Squared Error (MSE) as the evaluation metric for this task. To reduce the impact of extreme values, we calculate the MSE after applying the logarithmic transformation (log1p) to the true and predicted values.
Sara Entailment: In this task, given a statute, a fact, and an assertion, LLMs are required to determine if the assertion is "entailed" by the fact and statute. The dataset in this task is also from SARA, which tests the ability to reason about summaries of facts and statutes, in the context of US federal tax law.
Supreme Court Assessment of Legal Reasoning: In this task, the model must select, from a set of candidates, the holding statement that best answers a specific legal question. Each question represents an issue reviewed in a particular Supreme Court case, and the model must identify the holding statement that most accurately addresses it. This task is designed to assess legal reasoning by emphasizing the understanding of legal language over rote memorization of legal knowledge.
Legal Argument Reasoning: This task involves selecting the appropriate subsequent statement from multiple choices within a sequence of legal arguments presented during Court proceedings, based on the case facts. The input consists of a case description, a specific argument related to the case, and several potential candidate arguments. The objective is to determine which candidate argument logically continues the given argument.
3.2 Models
To evaluate the legal reasoning capabilities of LLMs, we test a range of models developed by different companies, including both general-purpose LLMs and those incorporating test-time scaling.
Table 3: LLMs used for legal reasoning evaluation
Category Model Name Source
General LLMs GPT-4o OpenAI
Llama-3.1-405B Meta
Qwen2.5-72B-Instruct Alibaba
Deekseek-V3 DeepSeek
Test Time Scaling LLMs DeepSeek-R1 DeepSeek
OpenAI-o1-preview OpenAI
OpenAI-o1-mini OpenAI
GLM-zero-preview Zhipu
QwQ-32B-Preview Alibaba
Table 3 summarizes the LLMs employed in our study. For general-purpose LLMs, we employ OpenAI’s GPT-4o, DeepSeek-V3, Meta’s Llama-3.1-405B, and Alibaba’s Qwen2.5-72B-Instruct. Meanwhile, for test-time scaling LLMs, we adopt DeepSeek-R1, OpenAI’s o1-preview, and o1-mini, alongside Zhipu’s GLM-zero-preview and Alibaba’s QwQ-32B-Preview.
3.3 Experiment Setting
During evaluation, we retrieve LLMs responses via API requests. Tailored prompts are designed for each task, ensuring a clear structure for the expected outputs. Additionally, API request parameters are adjusted based on the specific LLMs. Detailed task prompts can be found in Appendix A.
4 Experiment Results
In this section, we detail and analyze the performance of LLMs on legal tasks across both Chinese and English contexts.
4.1 Chinese Legal Task Results
Table 4: The result of Chinese legal task. The best score is bold, and the second-best score is underlined.
[…]
General LLMs
GPT-4o 59.40% 38.00% 65.00% 61.08% 61.79% 82.30% 82.33% 34.71% 65.99% 26.71% 41.33%
Llama3.1-405B 84.91% 35.86% 67.23% 49.26% 57.02% 84.94% 88.22% 34.25% 56.44% 20.03% 42.00%
Qwen2.5-72B-Instruct 80.34% 54.50% 76.50% 64.39% 61.41% 89.22% 84.97% 39.17% 60.50% 35.71% 50.67%
DeepSeek-V3 88.03% 45.00% 77.03% 63.94% 61.97% 87.67% 84.43% 40.00% 58.88% 36.05% 53.67%
Test Time Scaling LLMs
DeepSeek-R1 90.54% 71.67% 78.00% 68.48% 67.15% 91.71% 88.23% 42.80% 57.79% 55.91% 75.00%
o1-preview 90.13% 46.39% 76.63% 66.71% 63.05% 90.76% 74.07% 45.43% 66.83% 22.77% 55.03%
o1-mini 86.32% 27.00% 59.63% 42.59% 39.47% 84.02% 85.33% 39.78% 52.84% 12.62% 25.93%
GLM-zero-preview 72.22% 48.50% 71.98% 55.27% 55.82% 85.85% 78.56% 28.93% 57.00% 35.41% 48.67%
QwQ-32B-Preview 78.97% 56.00% 73.98% 60.70% 67.04% 87.04% 83.82% 27.75% 56.5% 34.93% 62.00%
As shown in table 4, DeepSeek-R1 demonstrates strong performance in Chinese tasks. However, all LLMs perform poorly on the CFE task, indicating its challenging nature. This may be due to the complex reasoning required, as the CFE task demands that models understand legal cases and infer the correct issue from a set of disputes. Moreover, the legal judgment prediction task for cases with multiple defendants proves more difficult for LLMs than for those with a single defendant. In contrast, the multi-defendant charge prediction task is relatively straightforward for LLMs, with DeepSeek-R1 achieving an accuracy of 91.71%—the highest among all tasks. Furthermore, the results indicate that most LLMs underperform on the AR task, suggesting that these models still struggle with recalling fine-grained legal knowledge. In conclusion, LLMs have difficulty with tasks that require complex reasoning, highlighting a key area for future improvement.
4.2 English Legal Task Results
Table 5: The result of English legal task. The best score is bold, and the second-best score is underlined.
[…]
General LLMs
GPT-4o 83.64% 82.41% 80.48% 1.207330476 87.87% 84.30% 81.73%
Llama3.1-405B 90.91% 61.11% 70.89% 7.71700315 80.88% 69.59% 83.84%
Qwen2-72B-Instruct 87.27% 82.41% 70.89% 4.806589399 85.29% 77.19% 77.89%
DeepSeek-V3 90.91% 77.78% 75.00% 2.307866126 83.09% 77.19% 85.00%
Test Time Scaling LLMs
DeepSeek-R1 92.73% 78.70% 77.05% 0.246035591 91.79% 85.28% 88.60%
o1-preview 96.36% 86.11% 79.79% 1.092967177 91.18% 86.98% 86.24%
o1-mini 87.27% 61.11% 66.78% 1.376229363 89.34% 73.53% 66.50%
GLM-zero-preview 83.64% 57.41% 65.41% 7.785547624 90.77% 70.76% 78.50%
QwQ-32B-Preview 78.18% 59.26% 64.73% 3.303448944 71.32% 73.41% 81.00%
As shown in table 5, the LLMs generally perform better on English reasoning tasks compared to Chinese ones. Furthermore, the performance of DeepSeek-R1 is comparable to o1-preview, with each achieving the best score three times. Additionally, the LRC and Sara_E tasks appear to be relatively easy for the LLMs, with the o1-preview achieving 96.36% and 91.18%, respectively. However, the NYSJE task seems to pose a challenge for the LLMs, with the highest accuracy being 80.48%. Moreover, most LLMs models struggle with the Sara_N task, except for DeepSeek-R1. These results suggest that most LLMs may currently lack the necessary capability to handle tasks related to judicial ethics and complex tax calculations.
5 Analysis
In this section, we further explore the performance of LLMs in two key legal reasoning tasks under both Chinese and English settings.
5.1 LLMs Performance in LJP Task
In this section, we evaluate the performance of LLMs on single-defendant and multi-defendant LJP tasks across three datasets: CAIL2018, CMDL, and MultiLJP. The primary difference between these tasks is the number of defendants involved. Additionally, we detail the results for each subtask within these datasets.
Table 6: The CAIL2018 subtask result. The best score is bold, and the second-best score is underlined.
[…]
General LLMs
GPT-4o 65.37% 90.67% 77.44% 37.33%
Llama3.1-405B 67.23% 86.00% 83.00% 41.33%
Qwen2-72B-Instruct 76.50% 93.44% 94.44% 50.33%
DeepSeek-V3 77.03% 94.44% 94.78% 50.67%
Test Time Scaling LLMs
DeepSeek-R1 78.00% 95.00% 95.67% 52.00%
o1-preview 76.63% 94.33% 96.67% 48.33%
o1-mini 59.63% 90.33% 62.67% 34.33%
GLM-zero-preview 71.98% 87.11% 93.28% 44.67%
QwQ-32B-Preview 73.98% 90.00% 92.17% 48.33%
Table 7: The CMDL subtask results. The best score is bold, and the second-best score is underlined.
[…]
General LLMs
GPT-4o 61.08% 85.57% 81.38% 27.50%
Llama3.1-405B 49.26% 77.86% 58.47% 20.91%
Qwen2-72B-Instruct 64.39% 87.57% 87.15% 29.92%
DeepSeek-V3 63.94% 88.05% 83.69% 31.03%
Test Time Scaling LLMs
DeepSeek-R1 68.48% 92.32% 91.19% 33.58%
o1-preview 66.71% 91.00% 91.29% 30.07%
o1-mini 42.59% 82.22% 30.98% 21.56%
GLM-zero-preview 55.27% 78.97% 73.07% 24.14%
QwQ-32B-Preview 60.70% 79.82% 82.66% 29.90%
Table 8: The MultiLJP subtask results. The best score is bold, and the second-best score is underlined.
[…]
General LLMs
GPT-4o 61.79% 84.72% 90.45% 23.1%
Llama3.1-405B 57.02% 74.13% 81.4% 25.9%
Qwen2-72B-Instruct 61.41% 83.82% 87.78% 24.84%
DeepSeek-V3 61.97% 81.95% 86.61% 28.51%
Test Time Scaling LLMs
DeepSeek-R1 67.15% 85.46% 92.15% 34.65%
o1-preview 63.05% 84.11% 89.14% 27.68%
o1-mini 39.47% 70.07% 36.73% 18.56%
GLM-zero-preview 55.82% 80.03% 68.62% 28.07%
QwQ-32B-Preview 67.04% 85.66% 89.6% 36.17%
From table 6,7 and 8 above, We can observe that the performance of LLMs on the CAIL2018 dataset is superior to that on CMDL and MultiLJP across all subtasks. This may be attributed to the relative simplicity of single-defendant cases, which facilitates comprehension and judgment prediction by the models. However, the o1-mini model performs worst in the article prediction task, with F1 scores of 30.98%, 36.73%, and 62.67% respectively. Additionally, the DeepSeek-R1 model demonstrates superior legal reasoning capabilities compared to the other LLMs in the LJP task.
5.2 LLMs Performance in Scalr Task
In this section, we further analyze the performance of LLMs in the scalr task. This task requires LLMs to comprehend legal language, rather than simply memorizing specific legal knowledge. Similarly to the LegalBench setting, we randomly sample 30% of the full dataset (comprising 172 questions) to serve as our test set. Following the results from LegalBench  [^4] https://hazyresearch.stanford.edu/legalbench/tasks/scalr.html, we compare the performance of some typical LLMs on this task.
Table 9: Performance comparison of different models in scalr task. The best score is bold, and the second-best score is underlined.
[…]
Human 86.00%
DeepSeek-R1 85.28%
o1-preview 86.98%
o1-mini 73.53%
GPT-4o 84.30%
gpt-4-0314 81.40%
gpt-3.5-turbo-0301 64.0%
As shown in table 9, the o1-preview model outperforms human performance, with DeepSeek-R1 ranking second. The latest OpenAI models exceed gpt-3.5-turbo-0301, and our evaluation reveals that recent LLMs also surpass gpt-3.5-turbo-0301, underscoring significant advancements in legal reasoning capabilities.
6 Error Analysis
Based on the experimental results in Section 4, DeepSeek-R1 outperforms other LLMs in most Chinese legal tasks and achieves performance comparable to o1-preview in English tasks. Consequently, we perform an error analysis of R1’s performance on several representative tasks. For Chinese legal tasks, we examine the IAPE, CFE, LJP, and AR tasks; for English legal tasks, we analyze the CPC and NYSJE tasks.
Figure 2: R1 error types across typical tasks.
6.1 IAPE Error Analysis
In the IAPE task, we randomly select 30 incorrect cases and identify two main error types: inconsistent objects and incomplete rebuttals. The inconsistent objects occurs when the model’s prediction doesn’t address the same legal subject as the given argument. The incomplete rebuttals happens when the prediction is not the best choice compared to the correct one. As shown in panel (a) of figure 2, 93% involved inconsistencies in the subject of discussion, while 7% had incomplete rebuttals. After analyzing the thinking process of R1, we conclude the reasons as follows:
(1) In the thinking process of R1, it fails to grasp the core subject in the plaintiff’s argument, making it hard to identify the correct iterative pair.
(2) The R1 fails to identify the argument that directly interacts with the plaintiff, instead selecting a response that comes after a lengthy reasoning process. This excessive reasoning leads to an incorrect answer.
We provide two examples of R1’s flawed thinking in the IAPE task, as shown in table 10.
Table 10: Examples of R1’s flawed thinking in the IAPE task.
Type Example
Missing Plaintiff’s Core Argument I need to identify which defense argument directly responds to or interacts with the plaintiff’s compensation claim. The plaintiff is seeking joint compensation, so I need to look for any defense that directly addresses the liability for compensation or the amount. bc_5 Insurance Company has made it clear that it will assume responsibility within the scope of compulsory insurance, which directly corresponds to the compensation amount and falls within the scope of the insurance company’s liability. (The core subject is compensation.)
Flawed and Excessive Reasoning Dispute over Liability Allocation: The plaintiff asserts through evidence that Sha should bear full responsibility, whereas bc_2 proposes that Deng xx should bear secondary responsibility. If Deng xx’s responsibility holds true, the scope of Sha’s responsibility will inevitably be curtailed (e.g., primary responsibility rather than full responsibility), directly affecting whether he constitutes a traffic accident crime (since this crime typically requires the perpetrator to bear major or full responsibility). Confrontation on Conviction and Sentencing: The plaintiff demands a heavier punishment for Sha, while bc_2 argues that Deng xx should not be deemed criminal and should be given a lighter punishment based on circumstances such as surrender and compensation. If Deng xx’s responsibility is acknowledged, it might undermine the foundation for Sha’s conviction and subsequently weaken the rationality of imposing a more severe punishment. Therefore, bc_2 is the sole argument that forms a substantial interaction with the plaintiff. By introducing the responsibility sharing of another defendant, it directly challenges the plaintiff’s core assertion regarding Sha’s full responsibility and the establishment of criminal liability. (A lengthy overthinking reasoning process)
6.2 CFE Error Analysis
In this task, we also randomly select 30 incorrect cases and category the errors into four types based on the number of disputes incorrectly extracted: complete deviation, major deviation, partial deviation, and minor deviation. In the 30 examples we randomly analyze, 67% exhibit complete deviation, while 13% show minor deviations. The remaining two categories account for 10%. We analyze the thinking process of R1 in error cases, concluding the reasons as follows:
(1) R1 lacks sufficient legal knowledge in certain areas, such as the concept of "an equal division of common property", which involves two key aspects: property division and property nature determination. However, R1 fails to identify the determination of property nature. Similarly, the issue of "arbitration requests for double wages exceeding the statute of limitations" involves two components: whether double wages are due and the statute of limitations for arbitration.
(2) R1 sometimes struggles with understanding more complex cases or the meaning of some disputed issues, which hinders accurate identification of the dispute issue.
(3) Some disputes, such as the determination of loss compensation amount and the determination of payment amount, are confusing to R1, making it difficult to identify the correct one.
The table 11 below provides examples illustrating these reasons.
Table 11: Examples of R1’s flawed thinking in the CFE task.
Type Example
Lacking Sufficient Legal Knowledge Additionally, the division of vehicles and real estate in property division may involve "real estate division" 118, but index 45 has already covered the overall property division. Moreover, the defendant’s request for an equal division of common property also falls within the scope of property division. There may also be "determination of property nature" 44, but there is no mention of a dispute over the nature of the property in the case, only the issue of division.
Struggling to Understand Disputed Issues The second part is the issue of advance payment for medical expenses, amounting to 7,900 yuan, which needs to be dealt with in this case. This involves the determination of medical expenses, which falls under "medical and nursing expense determination," corresponding to index 14. However, it might also involve the determination of compensation items or amounts. Looking more closely at the dictionary, there are "compensation item determination" (6) and "loss compensation amount determination" (7), but since medical expenses are specific medical and nursing costs, the more direct reference is index 14, "medical and nursing expense determination."
Confusing Disputed Issues The core dispute in the case facts lies in the appellant’s claim that the loan in question is a fictitious debt and believes that even if the loan is real, it belongs to PER’s personal debt rather than the joint debt of the couple. The appellee, on the other hand, claims that the loan is real and valid. According to the dispute focus dictionary, "determination of joint debt of husband and wife" (index 54) directly corresponds to the dispute between the two parties over whether the debt belongs to the joint debt during the marriage relationship, which is the most fitting dispute focus. (The ground truth is Joint Debt Claim of Husband and Wife / Common Debt during the Existence of Marriage Relationship. )
6.3 LJP Error Analysis
In the error analysis of the LJP task, we analyze the sentence prediction task of CAIL2018, CMDL, and MultiLJP. We also randomly select 30 wrong cases and divide the error types into 2 types, underestimation of the sentence and overestimation of the sentence.
As shown in figure2 (c), (d), and (e), in the CAIL2018, 63% of the cases involve an overestimation of the punishment, while 37% reflect an underestimation. In the MultiLJP, we observe that 60% of the incorrect cases involve an overestimation of the punishment, and 40% show an underestimation of the sentence. In the CMDL, 53% of the cases involve an underestimation of the sentence, while 47% reflect an overestimation of the punishment.
We conclude the reasons as follows:
(1) There exist some hallucinations in the thinking process, including highlighting or ignoring the sentencing factors and confusing the crimes. When R1 predicts the wrong charges of cases, the basic sentence is away from the ground truth, resulting in the wrong sentence.
(2) There are some outdated legal provisions in the training data, R1 predicts the sentence based on these articles, resulting the wrong sentence of defendants, such as the articles of operating a gambling house.
(3) We find that R1 predicts the sentence of the current case based on judgment of the similar cases in its training data. Therefore, when it selects the incorrectly similar case, the sentence is away from the ground truth.
Table 12 presents examples illustrating the three reasons.
Table 12: Examples of R1’s flawed thinking in the LJP task.
Type Example
Confusing the Crimes Based on provided charge categories, possible charges include contract fraud and fraud. The defendants obtained vehicles through a leasing agreement, then fabricated IOUs and sold the vehicles with the intent to illegally possess them, which meets the constituent elements of contract fraud. Contract fraud involves deceiving parties to obtain property during the signing or performance of contracts. Here, fraudulent acts were committed during the leasing process. Therefore, the charge should be contract fraud, corresponding to Article 224 of the Criminal Law.
Outdated Legal Provision Next, determine the sentence. According to Article 303 of the Criminal Law, the crime of operating a gambling establishment generally carries a sentence of up to 3 years of imprisonment, criminal detention, or controlled release, plus fines; for severe cases, 3 to 10 years. In this case, their illegal gains exceed 50,000 yuan, constituting a "relatively large amount" but likely not qualifying as a "severe circumstance".
Based on Similar Case Judgment However, considering compensation and victim reconciliation, the sentence may be moderately reduced. In practice, the combined sentence might range from 1 year and 6 months to 2 years. Based on similar precedent cases, the sentence could be set at 1 year and 6 months of fixed-term imprisonment.
6.4 AR Error Analysis
In the AR task, we also randomly select 30 incorrect cases from R1 and analyze the error types. We summarize four error types: confusion between articles, where the model incorrectly substitutes content from one legal article for another; content fabrication, where R1 generates an entirely fictional article that doesn’t exist in the actual legal corpus; incomplete recitation, where the model omits essential portions of the legal article; and outdated references, where R1 recites previous versions of articles that have since been amended or updated. As shown in figure 2 (f), we observe that in 73% of cases, the model creates confusion between articles, and 17% of cases involve fabricated articles. Besides, 7% of cases involve incomplete recitation, and 3% of cases are outdated references. Table 13 illustrates examples of these error types. observe
Table 13: Examples of R1’s flawed thinking in the AR task.
Error Types Example
Confusion between Articles Ground Truth: Article 17 of the Law on the Protection of Women’s Rights and Interests: The relevant departments should listen to and adopt criticisms or feasible suggestions regarding the safeguarding of women’s rights and interests.Prediction:Article 15 of the Law on the Protection of Women’s Rights and Interests: The state actively cultivates and selects female cadres, with emphasis on cultivating and selecting female cadres from ethnic minorities.
Content Fabrication Ground Truth: Article 51 of the Veterans Security Law: Veterans who meet the conditions for preferential housing arrangement shall be provided with a combination of market purchase and centralized construction by the military and local governments, which shall be planned and implemented scientifically by the people’s government of the resettlement area. Prediction: If the unit that receives and resettles demobilized military personnel violates the provisions of these regulations and falls under one of the following circumstances, the competent department of veterans’ affairs of the local people’s government shall order it to make corrections within a time limit.
Outdated References Ground Truth(Revised on February 24, 2017): Article 20 of the Red Cross Society Law: The Red Cross shall, in accordance with the law, accept donations of funds and materials from natural persons, legal persons, and other organizations, and issue public welfare donation receipts that are uniformly supervised (printed) by the financial department to the donors. Prediction(Amended on August 27, 2009): The sources and uses of the Red Cross funds shall be reported to the Red Cross Council every year and be subject to inspection and supervision by the People’s Government.
6.5 CPC Error Analysis
In the CPC task, we collect all the incorrect cases of this task from R1 and analyze the error types. As shown in figure 1 (g), the R1 model is more likely to misclassify the case labeled as "No" to "Yes" (57%) than the reverse (43%). We further analyze the reasons behind this:
(1) We find that there are factual hallucinations about the content of citation during the thinking process of R1. In addition, when the model is unclear about the details of the case, hallucinations will also occur, resulting in incorrect judgments.
(2) In this task, correctly understanding the citation is crucial for providing an accurate answer. Although R1 has access to the full case details, a deviation in understanding the case also leads to the wrong conclusion.
Table 14 presents examples illustrating the two reasons.
Table 14: Examples of R1’s flawed thinking in the CPC task.
Type Example
Citation Factual Inaccuracies Assuming that Lovern v. Edwards does discuss the reasonable jury standard in the context of summary judgment, the answer would be Yes. If not, then No. Since the user is asking for a Yes/No answer, I need to be certain. Without access to the full case details, I might rely on the common knowledge that Lovern v. Edwards is cited for summary judgment standards, which include the reasonable jury test. Therefore, the answer is likely Yes.
Misunderstanding the Citation The citation to Gunn v. Minton does not directly support the text. While Gunn addresses federal question jurisdiction, it focuses on whether a state-law claim "arises under" federal law when it implicates a substantial federal issue—not whether federal law creating the cause of action establishes jurisdiction. The text reflects the general rule from cases like Merrell Dow Pharmaceuticals Inc. v. Thompson, not Gunn.
6.6 NYSJE Error Analysis
We also collect all the incorrect cases of the NYSJE task from R1 and analyze the error types. As shown in figure 1 (h), we observe that False Positives and False Negatives account for nearly half of the incorrect cases. This task requires LLMs to determine whether a question violates judicial ethics in the New York State Unified Court System. According to the results in the table 5, R1 achieves an accuracy of 77.05%, while the best score is 80.48%, indicating that the task is challenging for R1. We further analyze the reasons behind this.
(1) We find that there are factual hallucinations in ethical guidelines that R1 recalls, which mislead R1 do the wrong conclusion.
(2) In the thinking process of R1, there is an overthinking phenomenon. For example, it adds irrelevant details to the context of the question, which results in incorrect output.
Table 15 provides examples illustrating the two reasons.
Table 15: Examples of R1’s flawed thinking in the NYSJE task.
Type Example
Ethical Guidelines Factual Hallucinations First, I need to recall the judicial ethics rules. Judges are supposed to follow the law, right? Even if they disagree with a law, they have to uphold it unless it’s been declared unconstitutional by a higher court. So if a judge believes a statute is unconstitutional, can they just ignore it? Probably not.
Adding Irrelevant Details But wait, sometimes there’s a rule that judges shouldn’t appear before other judges in the same county if it could create an appearance of favoritism. However, if both are part-time and their courts are separate, maybe it’s okay. The key is whether the Committee has a specific opinion on this.
7 Conclusion
This preliminary study evaluates the legal reasoning abilities of LLMs across 18 existing Chinese and English legal reasoning datasets using the latest LLMs models, including the TTS LLMs. We present our initial results and analysis, which provide evidence that the legal reasoning capabilities of TTS LLMs are still underdeveloped. The findings highlight that LLMs continue to struggle with both understanding legal terminology and executing complex legal reasoning. Future development of legal LLMs should prioritize improving these areas to enhance their overall legal reasoning performance.
8 Limitations
In this paper, we evaluate the legal reasoning abilities of LLMs using a collection of Chinese and English legal reasoning tasks. However, the tasks we gathered may not cover the full spectrum of legal reasoning scenarios. In the future, we plan to incorporate additional LLMs and expand the tasks to include subjective legal reasoning tasks. These tasks require applying legal provisions, similar cases, and principles through logical analysis of given facts and rules to reach conclusions. Additionally, defining appropriate evaluation metrics for this task is challenging. Therefore, we will explore and develop suitable metrics to better assess the performance of LLMs in these subjective legal reasoning tasks.
9 Reference
Appendix A Appendix
In this section, we present the instructions for LLMs in the evaluation of the Chinese and English legal tasks.
Figure 3: The prompt for LC dataset.
Figure 4: The prompt for LMHR dataset.
Figure 5: The prompt for CAIL2018 dataset.
Figure 6: The Chinese prompt for CAIL2018 dataset.
Figure 7: The prompt for CMDL dataset.
Figure 8: The Chinese prompt for CMDL dataset.
Figure 9: The prompt for MultiLJP dataset.
Figure 10: The prompt for MUD dataset.
Figure 11: The prompt for MSLRC dataset.
Figure 12: The prompt for CFE dataset.
Figure 13: The prompt for ArgMine dataset.
Figure 14: The prompt for AR dataset.
Figure 15: The prompt for JE dataset.
Figure 16: The prompt for LRC dataset.
Figure 17: The prompt for CPC dataset.
Figure 18: The prompt for NYSJE dataset.
Figure 19: The prompt for SARA_N dataset.
Figure 20: The prompt for SARA_E dataset.
Figure 21: The prompt for Scalr dataset.
Figure 22: The prompt for LAR dataset.